Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report page-20130130

Observers: Nancy Elias-Rosa, Antonia Morales-Garoffolo
Data reduction team: Sandra Benitez, Stefan Taubenberger

Log

Observations with EFOSC2. Cirrus from 05:30 UT on. In the afternoon there was a 6.8 earthquake and some softer replicas later on. 

Calibration data done in the afternoon.

UT 23:52   UBVRi skyflats

UT 00:29   Focus+Image Analysis

UT 00:45   SPECSTD 1745a         airmass=1.34    g11,13,16 sl 1.0

--DIMM SEEING 0.75"--

UT 00:59 LSQ12hot       We tried pointing to the object but there was some problem finding a guide and weren't sure if it was a pointing problem or something else so decided to go on to next object that was critical-critical (EFOSC2 doesn't like LSQ12hot)

UT 01:19   LSQ12heq              airmass=1.09    g13 sl 1.0

UT 04:00   OGLE-2012-SN-06       airmass=1.80     UBVRi

--DIMM SEEING 0.61"--

UT 04:43   OGLE-2013-SN-012      airmass=1.52    g13 sl1.0   **CLASSIFICATION Not enough flux.

UT 05:09   SPECSTD GD71          airmass=2.42    g11, 13, 16 sl 1.0

UT 05:20   Technical problems with EFOSC2

UT 05:38   OGLE-2013-SN-012      airmass=1.71    g13 sl1.0   **CLASSIFICATION               

--DIMM SEEING 1.16"--

UT 06:10   SN 2012hr             airmass=1.45    gri

UT 06:20   SN 2013K              airmass=2.16    g11,16 sl1.0 we increased exposure time of last spectrum because of the clouds

UT 09:13   SPECSTD GD153         airmass=1.60    g11,13,16 sl 1.0

UT 09:31   gri Skyflats

UT 09:52   END of the night       







Comments